### v2

# Accelerated decomposition of bistochastic kernel matrices by low rank approximation

Abstract
We develop an accelerated algorithm for computing an approximate eigenvalue decomposition of bistochastic normalized kernel matrices. Our approach constructs a low rank approximation of the original kernel matrix by the pivoted partial Cholesky algorithm and uses it to compute an approximate decomposition of its bistochastic normalization without requiring the formation of the full kernel matrix. The cost of the proposed algorithm depends linearly on the size of the employed training dataset and quadratically on the rank of the low rank approximation, offering a significant cost reduction compared to the naive approach. We apply the proposed algorithm to the kernel based extraction of spatiotemporal patterns from chaotic dynamics, demonstrating its accuracy while also comparing it with an alternative algorithm consisting of subsampling and Nyström extension.
1 Introduction
Data matrices of large size often arise in the application of data driven computational methods to various domains. An example is given by kernel matrices, whose entries are determined by the evaluation of a kernel function on a set of data points. Large kernel matrices arise in kernel based methods such as support vector machines, applied to tasks such as clustering and regression [41, 10]. Kernel methods enable the use of linear computational methods to capture nonlinear relationships in the underlying data. In addition, they facilitate the application of functional analytic concepts to datasets that belong to sets without additional mathematical structure, such as the structure of vector spaces or manifolds.
Despite their large size, data matrices often have a relatively low approximate rank [45]. Their underlying low rank structure manifests itself in fast spectral decay and can be exploited to compute low rank approximations. In turn, these approximations can be used to accelerate downstream computations with only a moderate loss in accuracy. However, explicitly computing the eigenvalue or singular value decomposition of a data matrix to construct its low rank approximation is often unfeasible in practice. This creates the need for efficient and effective low rank approximation methods that can scale to matrices of large size [28, 36, 37, 11, 19].
In this work we are specifically interested in building low rank approximations of bistochastic kernel matrices and using them to compute their eigenvalue decomposition with reduced cost, thereby enabling us to apply bistochastic kernel based methods to datasets of large size. In the remainder of Section 1 we briefly discuss kernel matrices, a class of methods for their low rank approximation, and the contributions of the present work. In Section 2 we outline previous related work that provides the motivation for the material to follow. In Section 3 we present two algorithms for accelerating the computation of the eigenvalue decomposition of bistochastic kernel matrices by low rank approximation. In Section 4 we apply the two algorithms to extract spatiotemporal patterns from simulation data of chaotic dynamics, followed by a brief conclusion in Section 5. The code used to generate the numerical results and plots of this work can be accessed online [^1] https://github.com/cval26/kernel_evd.
1.1 Kernel matrices
We consider a continuous, bounded and positive definite kernel function $k\colon X\times X\to\mathbb{R}$ taking nonnegative values, where the state space $X=\mathbb{R}^{d}$, $d\in\mathbb{N}$ is equipped with a probability measure $\mu$ with compact support. Given a collection of $N\in\mathbb{N}$ state samples $X_{N}=\{x_{i}\}_{i=0}^{N-1}\subset X$ we build the kernel matrix $\bm{K}\in\mathbb{R}^{N\times N}$ whose entries $K_{ij}=k(x_{i},x_{j})$ correspond to evaluation of kernel function $k$ on the given states. By construction, the kernel matrix $\bm{K}$ is positive definite, meaning that it is symmetric and that $\underline{y}^{\top}\bm{K}\underline{y}>0$ for every nonzero $\underline{y}\in\mathbb{R}^{N}$. As a consequence, it has a well defined eigenvalue decomposition with real positive eigenvalues and orthogonal eigenvectors. In what follows, we denote by $\mu_{N}=\sum_{i=0}^{N-1}\delta_{x_{i}}/N$ the empirical sampling measure associated with the state samples $X_{N}$. In applications we only have access to the finite collection of samples $X_{N}$ and so we use the sampling measure $\mu_{N}$ to approximate $\mu$, assuming an appropriate form of weak convergence as $N\to\infty$.
It is often desirable in applications to normalize the kernel function $k$ while maintaining its symmetry. One way to achieve that is to define the normalized kernel $\ell\colon X\times X\to\mathbb{R}$
$\ell(x,y)=\frac{k(x,y)}{\sqrt{d(x)}\sqrt{d(y)}}$
with normalization function $d\colon X\to\mathbb{R}$
$d(x)=\int_{X}k(x,y)d\mu(y)$
assuming that $d(x)>0$ for all $x\in X$. Although normalized, the integral $\int_{X}\ell(x,y)d\mu(y)$ is not necessarily equal to one for every $x\in X$, meaning that $\ell$ is generally not a stochastic (also called markovian) kernel. In the discrete setting, $\mu$ is replaced by $\mu_{N}$ and the above normalization procedure corresponds to forming the $N\times N$ normalized kernel matrix
$\bm{L}=\bm{D}^{-1/2}\bm{K}\bm{D}^{-1/2}$
with diagonal matrix $\bm{D}=\operatorname{diag}(\bm{K}\underline{1}_{N})$ holding the row sums of $\bm{K}$ in its diagonal and $\underline{1}_{N}\in\mathbb{R}^{N}$ denoting the unit vector. Normalized kernel matrices of this form are often employed in diffusion maps algorithms [13, 7, 8] and applications such as kernel spectral clustering (Section 2) [38, 47].
Normalizing the kernel $k$ in a way that turns it into a stochastic kernel while maintaining its symmetry can be achieved with the bistochastic normalization procedure developed in [12]. The bistochastic kernel function $p\colon X\times X\to\mathbb{R}$ is defined as
$p(x,y)=\int_{X}\frac{k(x,z)k(z,y)}{d(x)q(z)d(y)}\,d\mu(z)$
with positive functions $d:X\to\mathbb{R}$ and $q:X\to\mathbb{R}$
$d(x)=\int_{X}k(x,y)d\mu(y)\qquad q(x)=\int_{X}\frac{k(x,y)}{d(y)}d\mu(y).$
The analogous discrete procedure involves forming the $N\times N$ bistochastic kernel matrix
$\bm{P}=\bm{D}^{-1}\bm{K}\bm{Q}^{-1}\bm{K}\bm{D}^{-1}$
with diagonal matrices
$\bm{D}=\operatorname{diag}(\bm{K}\underline{1}_{N})\qquad\bm{Q}=\operatorname{diag}(\bm{K}\bm{D}^{-1}\underline{1}_{N}).$
Being stochastic, $\bm{P}$ can be interpreted as the transition probability matrix of a Markov chain of $N$ states. In addition, its eigenvalues are within the interval $0,1]\subset\mathbb{R}$, with its largest eigenvalue being equal to one and having a constant corresponding eigenvector.
Maintaining symmetry while normalizing a kernel matrix ensures that the resulting matrix has an eigenvalue decomposition. In the continuous setting, this means that the eigenfunctions of the associated kernel integral operator define an orthonormal basis of $L^{2}(X,\mu)$. We are going to use this fact to define the spatiotemporal patterns of a dynamical system in Section 4. As mentioned above, one motivating reason for working with a bistochastic kernel is that the associated kernel integral operator’s largest eigenvalue is equal to one and has a corresponding constant eigenfunction. This property is useful for proving convergence of Galerkin approximation schemes based on the obtained eigenfunctions [17]. Finally, ensuring that the eigenvalues of a symmetric matrix are within the interval $[0,1]$ has computational benefits as well, since it eliminates the possibility of unbounded growth of a vector under repeated applications of the matrix.
1.2 Partial Cholesky factorization
The Cholesky factorization of a positive semidefinite matrix $\bm{A}\in\mathbb{R}^{N\times N}$ is a decomposition of the form $\bm{A}=\bm{F}\bm{F}^{\top}$ with lower triangular Cholesky factor $\bm{F}\in\mathbb{R}^{N\times N}$ [44]. Because $\bm{A}$ is only assumed to be positive semidefinite, instead of positive definite, its Cholesky factorization is generally not unique. In this work we focus on the partial Cholesky factorization, which is an approximate decomposition
$\bm{A}\approx\tilde{\bm{A}}=\bm{F}\bm{F}^{\top}$
where the partial Cholesky factor $\bm{F}\in\mathbb{R}^{N\times r}$ need not be lower triangular anymore and $\operatorname{rank}{\tilde{\bm{A}}}\leq r$ with rank parameter $r<N$.
The partial Cholesky factorization can be used to compute low rank approximations of positive semidefinite matrices by the pivoted partial Cholesky algorithm [11, 19]. Provided a rank parameter $r$, the algorithm selects $r$ column indices (pivots) of the input matrix $\bm{A}$ and uses the corresponding $r$ columns to compute the approximate factorization $\tilde{\bm{A}}=\bm{F}\bm{F}^{\top}$. The approximation computed by the pivoted partial Cholesky algorithm for a given set of column indices $S=\{s_{0},\ldots,s_{r-1}\}$ is equal to the column Nyström approximation
$\bm{A}\approx\tilde{\bm{A}}=\bm{A}_{S}(\bm{A}_{S}^{S})^{+}\bm{A}_{S}^{\top}$
where $\bm{A}_{S}\in\mathbb{R}^{N\times r}$ denotes the submatrix of $\bm{A}$ formed by the columns indexed by the set $S$, $\bm{A}_{S}^{S}\in\mathbb{R}^{r\times r}$ the submatrix formed by the rows and columns indexed by $S$, and superscript ++ the Moore-Penrose pseudoinverse [30, 1, 11]. The above implies that the resulting approximation $\tilde{\bm{A}}$ computed by the pivoted partial Cholesky algorithm corresponds to a scaled projection of $\bm{A}$ onto the linear span of the $r$ selected columns, returned in factored form.
The column indices sampled by the pivoted partial Cholesky algorithm are selected iteratively based on the diagonal entries of the input matrix $\bm{A}$, with different sampling strategies leading to different variations of the algorithm and different degrees of accuracy of the resulting low rank approximation. The indices selected by the employed algorithm correspond to states identified as belonging to the most “important” states of the original dataset according to the used sampling strategy. Namely, every pivoted partial Cholesky algorithm is in effect also a sampling algorithm that can be used to subsample a given dataset in an informed manner.
In this work we employ the adaptive random sampling strategy developed in [11], using the relative trace norm error $\operatorname{tr}{(\bm{A}-\tilde{\bm{A}})}/\operatorname{tr}{\bm{A}}$ to measure the degree of accuracy of the resulting approximation. Overall, the asymptotic computational cost of computing a rank-$r$ partial Cholesky factorization of an $N\times N$ matrix $\bm{A}$ is $O(Nr^{2})$. Importantly, when $\bm{A}$ is a kernel matrix, the method requires the evaluation of the associated kernel function almost exclusively on the sampled columns, resulting in only $N(r+1)$ kernel evaluations.
For an $N\times N$ positive semidefinite matrix $\bm{A}$ and a rank parameter $r^{\prime}\in\mathbb{N}$ we denote by $\bm{A}]_{r^{\prime}}\in\mathbb{R}^{N\times N}$ a best rank-$r^{\prime}$ approximation of $\bm{A}$, which is generally not unique. As before, we denote by $\tilde{\bm{A}}\in\mathbb{R}^{N\times N}$ the rank-$r$ approximation produced by the randomly pivoted partial Cholesky algorithm with the sampling strategy of [11], for a rank parameter $r\geq r^{\prime}$. For fixed parameters $r^{\prime}$ and $\epsilon>0$ it is shown in [11] that the approximation $\tilde{\bm{A}}$ satisfies the error bound
$\mathbb{E}\,\frac{\operatorname{tr}{(\bm{A}-\tilde{\bm{A}})}}{\operatorname{tr}{\bm{A}}}\leq(1+\epsilon)\,\eta\qquad\eta=\operatorname{tr}{(\bm{A}-[\bm{A}]_{r^{\prime}})}/\operatorname{tr}{\bm{A}}$ (1.1)
for every $r$ such that
$r\geq\frac{r^{\prime}}{\epsilon}+r^{\prime}\log\frac{1}{\epsilon\eta}$
where the average is taken over all randomly sampled columns. Namely, when the number of columns $r$ of the partial Cholesky factorization satisfies the above inequality, then the relative trace norm error of the approximation is at most $\epsilon$ times greater than the best relative error for an approximation of lower rank $r^{\prime}\leq r$. We refer the reader to [11] for a precise statement of the proven error bounds and comparison to other sampling strategies. We will use the above approximation error bound to offer an explanation for some of the numerical results presented in Section 4.2.
1.3 Contributions
Our contributions can be summarized as follows.
  1. 1.
    Accelerated decomposition. Employing a rank-$r$ partial Cholesky factorization of an $N\times N$ kernel matrix, we develop an algorithm for the accelerated computation of the eigenvalue decomposition of its bistochastic normalized version. The accelerated algorithm has an asymptotic computational cost $O(Nr^{2})$ compared to the naive $O(N^{3})$, while requiring only $N(r+1)$ kernel evaluations. In addition, we compare the developed algorithm with an alternative approach based on subsampling and Nyström extension.
  2. 2.
    Application. We apply the two algorithms to the kernel based extraction of spatiotemporal patterns from chaotic dynamics, offering empirical evidence for their relative performance while also investigating their differences.
2 Related work
There is a large body of literature on the use of low rank approximation methods to accelerate the implementation of kernel methods and enable their application to large datasets. In this work we are interested in kernel methods that require the computation of the eigenvalue decomposition of a kernel matrix, with kernel spectral clustering being a prime example [38, 49, 47]. More specifically, we focus on two approximation methods, which we refer to as dilution and subsampling.
2.1 Dilution
We begin by considering the method of dilution, variations of which have been used in the past to accelerate the application of kernel spectral clustering [21, 20, 11]. We consider a kernel function $k\colon X\times X\to\mathbb{R}$ satisfying the assumptions listed in Section 1.1. Using a collection of $N\in\mathbb{N}$ state samples $X_{N}=\{x_{i}\}_{i=0}^{N-1}\subset X$ we build the corresponding kernel matrices $\bm{K}=[k(x_{i},x_{j})]_{i,j}\in\mathbb{R}^{N\times N}$ and
$\bm{L}=\bm{D}^{-1/2}\bm{K}\bm{D}^{-1/2}\in\mathbb{R}^{N\times N}$
with diagonal matrix $\bm{D}=\operatorname{diag}(\bm{K}\underline{1}_{N})$. Our goal is to compute the eigenvalue decomposition (EVD) of matrix $\bm{L}$, which is an operation of asymptotic cost $\mathcal{O}(N^{3})$. We operate under the assumption that solving the eigenvalue problem for $\bm{L}$ is computationally unfeasible due to the large number of samples $N$. For this reason, we are going to perform a low rank approximation of $\bm{L}$ and use it to reduce the cost of the eigenvalue problem
To that end we employ a pivoted partial Cholesky algorithm to build an approximation of matrix $\bm{K}$ that is of lower rank $r<N$
$\bm{K}\approx\tilde{\bm{K}}=\bm{F}\bm{F}^{\top}\in\mathbb{R}^{N\times N}$
with partial Cholesky factor $\bm{F}\in\mathbb{R}^{N\times r}$. To ensure that the resulting approximation of matrix $\bm{L}$ is properly normalized we build the corresponding diagonal matrix $\tilde{\bm{D}}=\operatorname{diag}(\tilde{\bm{K}}\underline{1}_{N})$ and form the normalized low rank approximation
$\bm{L}\approx\tilde{\bm{L}}=\tilde{\bm{D}}^{-1/2}\tilde{\bm{K}}\tilde{\bm{D}}^{-1/2}=(\tilde{\bm{D}}^{-1/2}\bm{F})(\tilde{\bm{D}}^{-1/2}\bm{F})^{\top}\in\mathbb{R}^{N\times N}.$
Our goal is to compute the EVD of matrix $\tilde{\bm{L}}$.
Notice that the above factorization of $\tilde{\bm{L}}$ is rank revealing, since the factor matrix $\tilde{\bm{D}}^{-1/2}\bm{F}$ is of reduced dimension $N\times r$. This means that we can compute the singular value decomposition (SVD) of $\tilde{\bm{D}}^{-1/2}\bm{F}$ with reduced cost $\mathcal{O}(Nr^{2})$ and use it to obtain the EVD of $\tilde{\bm{L}}$. More specifically, using the reduced SVD
$\tilde{\bm{D}}^{-1/2}\bm{F}=\bm{U}\bm{\Sigma}\bm{V}^{\top}$
the desired EVD of $\tilde{\bm{L}}$ reads
$\tilde{\bm{L}}=\bm{U}\bm{\Sigma}^{2}\bm{U}^{\top}$
with matrix $\bm{U}\in\mathbb{R}^{N\times r}$ holding the $r$ eigenvectors and diagonal matrix $\bm{\Sigma}^{2}\in\mathbb{R}^{r\times r}$ the corresponding eigenvalues, which act as respective approximations of the leading $r$ eigenvectors and eigenvalues of the original matrix $\bm{L}$. The approximation scheme outlined above can be carried out with an asymptotic computational cost $O(Nr^{2})$. Numerical results demonstrating its accuracy for kernel spectral clustering are presented in [11], where the authors employ their proposed adaptive random sampling strategy to build the partial Cholesky factor $\bm{F}$.
2.2 Subsampling
The method of subsampling has been applied to accelerate various kernel based methods, where it is usually referred to as Nyström extension or the Nyström method [48, 18, 3, 5, 4, 33, 35]. In this method, we identify a subset of the original dataset, build the associated kernel matrix, compute its eigenvectors and then extend them to the full dataset using the Nyström extension method.
Employing the pivoted partial Cholesky algorithm as a sampling algorithm, we extract a subset of $r\in\mathbb{N}$ state samples $\tilde{X}_{r}=\{\tilde{x}_{i}\}_{i=0}^{r-1}\subset X_{N}$ from the original dataset. Relying on those states only, we form the $r\times r$ kernel matrices $\tilde{\bm{K}}$ and $\tilde{\bm{L}}$ in the same way as earlier. Given its reduced dimensions, we compute the EVD of $\tilde{\bm{L}}$ directly
$\tilde{\bm{L}}=\tilde{\bm{\Phi}}\bm{\Lambda}\tilde{\bm{\Phi}}^{\top}$
with $\tilde{\Phi}$ holding its eigenvectors and $\Lambda$ its eigenvalues.
Each eigenvector (column) $\tilde{\underline{\phi}}_{i}$ of $\tilde{\bm{\Phi}}$ corresponding to a nonzero eigenvalue $\lambda_{i}$ can be thought of as holding the values of a continuous function $\tilde{\phi}_{i}:X\to\mathbb{R}$ on the set of points $\tilde{X}_{r}\subset X$. We can use the Nyström extension method to continuously extend each such eigenvector to hold the function’s values on the full dataset $X_{N}\subset X$. For every $x\in X_{N}$ we have
$\tilde{\phi}_{i}(x)=\frac{1}{\lambda_{i}}\int_{\tilde{X}_{r}}\ell(x,y)\tilde{\phi}_{i}(y)d\tilde{\mu}_{r}(y)=\frac{1}{r\lambda_{i}}\sum_{j=0}^{r-1}\ell(x,\tilde{x}_{j})\tilde{\phi}_{i}(\tilde{x}_{j})$
with $\tilde{\mu}_{r}=\sum_{i=0}^{r-1}\delta_{\tilde{x}_{i}}/r$ the empirical sampling measure associated with $\tilde{X}_{r}$. We collect the extended eigenvectors in matrix $\hat{\bm{\Phi}}\in\mathbb{R}^{N\times r}$ which is generally not orthogonal. To ensure orthogonality we compute the QR decomposition $\hat{\bm{\Phi}}=\bm{\Phi}\bm{R}$ where the orthogonal matrix $\bm{\Phi}\in\mathbb{R}^{N\times r}$ holds the approximate leading $r$ eigenvectors of matrix $\bm{L}$. The method’s overall asymptotic computational cost is $O(Nr^{2})$.
3 Bistochastic kernel approximation
In this section we extend the approximation schemes outlined in Section 2 to the case of bistochastic normalized kernel matrices (Section 1.1).
We consider again a kernel function $k\colon X\times X\to\mathbb{R}$ satisfying the assumptions listed in Section 1.1. We use a collection of $N\in\mathbb{N}$ state samples $X_{N}=\{x_{n}\}_{n=0}^{N-1}\subset X$ to build the kernel matrix $\bm{K}=[k(x_{i},x_{j})]_{i,j}\in\mathbb{R}^{N\times N}$ and its bistochastic normalization
$\bm{P}=\bm{D}^{-1}\bm{K}\bm{Q}^{-1}\bm{K}\bm{D}^{-1}\in\mathbb{R}^{N\times N}$
with $N\times N$ diagonal matrices $\bm{D}=\operatorname{diag}(\bm{K}\underline{1}_{N})$ and $\bm{Q}=\operatorname{diag}(\bm{K}\bm{D}^{-1}\underline{1}_{N})$. Our goal is to compute the EVD of the bistochastic matrix $\bm{P}$, which involves a cost of $\mathcal{O}(N^{3})$. Assuming again that $N$ is sufficiently large to render computing the EVD of $\bm{P}$ unfeasible, we are interested in using a partial Cholesky factorization of $\bm{K}$ to obtain an approximate EVD of $\bm{P}$ with reduced cost.
Using a pivoted partial Cholesky algorithm we compute a low rank approximation $\tilde{\bm{K}}$ of matrix $\bm{K}$ with rank $r<N$
$\bm{K}\approx\tilde{\bm{K}}=\bm{F}\bm{F}^{\top}$ (3.1)
with partial Cholesky factor $\bm{F}\in\mathbb{R}^{N\times r}$. Assuming that the normalization matrices are well defined (Section 3.3)
$\tilde{\bm{D}}=\operatorname{diag}(\tilde{\bm{K}}\underline{1}_{N})\qquad\tilde{\bm{Q}}=\operatorname{diag}(\tilde{\bm{K}}\tilde{\bm{D}}^{-1}\underline{1}_{N})$
we form the low rank approximation of $\bm{P}$
$\tilde{\bm{P}}=\tilde{\bm{D}}^{-1}\tilde{\bm{K}}\tilde{\bm{Q}}^{-1}\tilde{\bm{K}}\tilde{\bm{D}}^{-1}\in\mathbb{R}^{N\times N}.$
We are interested in computing the EVD of the approximate kernel matrix $\tilde{\bm{P}}$ by taking advantage of its lower rank to perform operations that depend primarily on rank parameter $r<N$. We note that $\tilde{\bm{P}}$ can be written in the factored form
$\tilde{\bm{P}}=\bm{G}\bm{G}^{\top}$
with factor $\bm{G}=\tilde{\bm{D}}^{-1}\tilde{\bm{K}}\tilde{\bm{Q}}^{-1/2}\in\mathbb{R}^{N\times N}$. Contrary to the normalized kernels considered in Section 2, the above factorization of kernel matrix $\tilde{\bm{P}}$ is not rank revealing, since $\bm{G}$ is of dimension $N\times N$. As a result, computing the SVD of $\bm{G}$ to arrive to the EVD of $\bm{G}\bm{G}^{\top}$ does not offer an immediate reduction in computational cost.
Algorithm 1 Dilution algorithm (Section 3.1).
 Input: state samples $X_{N}=\{x_{n}\}_{n=0}^{N-1}$ and approximation rank $r<N$
 Output: eigenvector matrix $\bm{\Phi}\in\mathbb{R}^{N\times r}$ and diagonal eigenvalues matrix $\bm{\Lambda}\in\mathbb{R}^{r\times r}$
$\bm{F},\,-\leftarrow\texttt{rpcholesky}\,(X_{N},r)$ partial Cholesky factor $\bm{F}$
$\tilde{\bm{D}}\leftarrow\texttt{diag}\,(\bm{F}\bm{F}^{\top}\underline{1}_{N}$)diagonal normalization matrix
assert $\texttt{diag}\,(\tilde{\bm{D}})>0$ ensure strict positivity of row sums
$\tilde{\bm{Q}}\leftarrow\texttt{diag}\,(\bm{F}\bm{F}^{\top}\tilde{\bm{D}}^{-1}\underline{1}_{N}$)diagonal normalization matrix
$\bm{V},\,\bm{\Sigma}\leftarrow\texttt{evd}\,(\bm{F}^{\top}\bm{F})$$r\times r$ EVD
$\bm{U}\leftarrow\bm{F}\bm{V}\bm{\Sigma}^{-1}$ left singular vectors of $\bm{F}$
$\bm{Q}_{1},\bm{R}_{1}\leftarrow\texttt{qr}\,(\tilde{\bm{D}}^{-1}\bm{U})$$N\times r$ reduced QR
$\bm{R}_{2}\leftarrow\texttt{qr}\,(\tilde{\bm{Q}}^{-1/2}\bm{U})$$N\times r$ reduced QR
$\bm{U}_{1},\bm{\Sigma}_{1},-\leftarrow\texttt{svd}\,(\bm{R}_{1}\bm{\Sigma}^{2}\bm{R}_{2}^{\top})$$r\times r$ SVD
$\bm{\Lambda}\leftarrow\bm{\Sigma}_{1}^{2}$ approximate eigenvalues
$\bm{\Phi}\leftarrow\bm{Q}_{1}\bm{U}_{1}$ approximate eigenvectors
$\bm{\Phi}(:,0)\leftarrow\underline{1}_{N}$(optional) set first column to 1
$\bm{\Phi},-\leftarrow\texttt{qr}\,(\bm{\Phi})$ (optional) re-orthonormalize eigenvectors
3.1 Dilution
We use the low rank approximation (3.1) to design an accelerated algorithm of reduced overall cost. To that end we begin by computing the EVD of $\bm{F}^{\top}\bm{F}\in\mathbb{R}^{r\times r}$
$\bm{F}^{\top}\bm{F}=\bm{V}\bm{\Sigma}^{2}\bm{V}^{\top}.$
The eigenvectors of matrix $\bm{F}^{\top}\bm{F}$ (columns of $\bm{V}$) are the right singular vectors of $\bm{F}$. Assuming that $\bm{F}^{\top}\bm{F}$ is of full rank we recover the left singular vectors of $\bm{F}$ by computing
$\bm{U}=\bm{F}\bm{V}\bm{\Sigma}^{-1}\in\mathbb{R}^{N\times r}.$
Using matrices $\bm{U}$ and $\bm{\Sigma}^{2}$ and the fact that $\tilde{\bm{K}}=\bm{F}\bm{F}^{\top}$ we write
$\bm{G}=\tilde{\bm{D}}^{-1}\bm{U}\bm{\Sigma}^{2}\bm{U}^{\top}\tilde{\bm{Q}}^{-1/2}=(\tilde{\bm{D}}^{-1}\bm{U})\bm{\Sigma}^{2}(\tilde{\bm{Q}}^{-1/2}\bm{U})^{\top}.$
Computing the reduced QR decompositions of the $N\times r$ matrices $\tilde{\bm{D}}^{-1}\bm{U}=\bm{Q}_{1}\bm{R}_{1}$ and $\tilde{\bm{Q}}^{-1/2}\bm{U}=\bm{Q}_{2}\bm{R}_{2}$ yields
$\bm{G}=\bm{Q}_{1}\bm{R}_{1}\bm{\Sigma}^{2}(\bm{Q}_{2}\bm{R}_{2})^{\top}=\bm{Q}_{1}(\bm{R}_{1}\bm{\Sigma}^{2}\bm{R}_{2}^{\top})\bm{Q}_{2}^{\top}.$
Finally, computing the SVD of the $r\times r$ square matrix $\bm{R}_{1}\bm{\Sigma}^{2}\bm{R}_{2}^{\top}=\bm{U}_{1}\bm{\Sigma}_{1}\bm{V}_{1}^{\top}$ leads to the desired factorization
$\bm{G}=(\bm{Q}_{1}\bm{U}_{1})\bm{\Sigma}_{1}(\bm{Q}_{2}\bm{V}_{1})^{\top}$ (3.2)
where the orthogonal matrix $\bm{Q}_{1}\bm{U}_{1}\in\mathbb{R}^{N\times r}$ holds the left singular vectors of $\bm{G}$ and the diagonal matrix $\bm{\Sigma}_{1}\in\mathbb{R}^{r\times r}$ the corresponding singular values. With those quantities known, the EVD of matrix $\tilde{\bm{P}}$ reads
$\tilde{\bm{P}}=\bm{G}\bm{G}^{\top}=(\bm{Q}_{1}\bm{U}_{1})\bm{\Sigma}_{1}^{2}(\bm{Q}_{1}\bm{U}_{1})^{\top}=\bm{\Phi}\bm{\Lambda}\bm{\Phi}^{\top}$ (3.3)
with $\bm{\Phi}=\bm{Q}_{1}\bm{U}_{1}\in\mathbb{R}^{N\times r}$ holding its eigenvectors and $\bm{\Lambda}=\bm{\Sigma}_{1}^{2}\in\mathbb{R}^{r\times r}$ its eigenvalues. Since $\tilde{\bm{P}}$ is bistochastic, we know that its leading eigenvector is a constant vector and its leading eigenvalue is equal to one. Taking advantage of that knowledge, one can optionally set the first column of $\bm{\Phi}$ to be the unit vector and perform another QR decomposition to re-orthonormalize its columns, potentially increasing the accuracy of the computed eigenvectors. Pseudocode for the method is given in Algorithm 1.
The dilution method involves an asymptotic computational cost $O(Nr^{2})$ for the rank-$r$ partial Cholesky factorization of $\bm{K}$, $O(r^{3})$ for the EVD of an $r\times r$ matrix, $O(r^{3})$ for the SVD of an $r\times r$ matrix, and $O(Nr^{2})$ for the QR decompositions and intermediate matrix products. Namely, the overall asymptotic cost is $O(Nr^{2})$. Additionally, the only computations that depend on the original size parameter $N$ are the QR decompositions and matrix products, which are operations with good parallel performance. To compute the QR decompositions efficiently one can use the classical Gram-Schmidt algorithm with reorthogonalization, which is an algorithm with good parallel performance that yields orthogonalization errors near machine precision [27].
Algorithm 2 Subsampling algorithm (Section 3.2).
 Input: state samples $X_{N}=\{x_{n}\}_{n=0}^{N-1}$ and approximation rank $r<N$
 Output: eigenvector matrix $\bm{\Phi}\in\mathbb{R}^{N\times r}$ and diagonal eigenvalues matrix $\bm{\Lambda}\in\mathbb{R}^{r\times r}$
$\tilde{X}_{r}\leftarrow\texttt{rpcholesky}\,(X_{N},r)$ subsampled dataset $\tilde{X}_{r}\subset X_{N}$
$\tilde{\bm{K}}\leftarrow\texttt{kernel}\,(\tilde{X}_{r})$kernel matrix
$\tilde{\bm{D}}\leftarrow\texttt{diag}\,(\tilde{\bm{K}}\underline{1}_{N}$)diagonal normalization matrix
$\tilde{\bm{Q}}\leftarrow\texttt{diag}\,(\tilde{\bm{K}}\tilde{\bm{D}}^{-1}\underline{1}_{N}$)diagonal normalization matrix
$\tilde{\bm{P}}\leftarrow\tilde{\bm{D}}^{-1}\tilde{\bm{K}}\tilde{\bm{Q}}^{-1}\tilde{\bm{K}}\tilde{\bm{D}}^{-1}$bistochastic normalized kernel matrix
[…]
$\hat{\bm{\Phi}}\leftarrow\texttt{extend}\,(\tilde{\bm{\Phi}},\tilde{X}_{r},X_{N})$Nyström extension
[…]
3.2 Subsampling
The method of subsampling can be applied to bistochastic normalized kernels following the same sequence of steps outlined in Section 2.2. Namely, we use the pivoted partial Cholesky algorithm to subsample the given dataset $X_{N}$, extracting a subset of $r$ samples $\tilde{X}_{r}\subset X_{N}$. Using only the extracted subset of samples we form the $r\times r$ kernel matrices $\tilde{\bm{K}}$ and $\tilde{\bm{P}}$ and compute the EVD of $\tilde{\bm{P}}$ directly.
The final step involves continuously extending the obtained eigenvectors to the full dataset $X_{N}$ using the Nyström extension method. In analogy to Section 2.2, for every $x\in X_{N}$ and each $\phi_{i}$ corresponding to a nonzero eigenvalue $\lambda_{i}$ we have the formula
$\phi_{i}(x)=\frac{1}{\lambda_{i}}\int_{\tilde{X}_{r}}p(x,y)\tilde{\phi}_{i}(y)d\tilde{\mu}_{r}(y)=\frac{1}{r\lambda_{i}}\sum_{j=0}^{r-1}p(x,\tilde{x}_{j})\phi_{i}(\tilde{x}_{j}).$
The desired eigenvectors are computed by performing the QR factorization of the matrix containing the extended eigenvectors. Pseudocode for the method is given in Algorithm 2.
The subsampling method involves an asymptotic computational cost $O(Nr^{2})$ for extracting the subset of $r$ state samples $\tilde{X}_{r}$ by the pivoted partial Cholesky algorithm, $O(r^{3})$ for the EVD of $\tilde{\bm{P}}$, $O(Nr^{3})$ for the Nyström extension and $O(Nr^{2})$ for the QR factorization. In total, the subsampling method’s asymptotic cost is $O(Nr^{3})$. Generally, the nonlocal extension formula used for the Nyström extension carries a cost $O(Nr^{2})$. However, in our case the evaluation of kernel $p$ is also nonlocal, yielding the worse scaling $O(Nr^{3})$. Despite its increased asymptotic cost, the Nyström extension is an inherently parallel computation and can thus be accelerated significantly.
3.3 Discussion
When applying the dilution method, the low rank approximation $\tilde{\bm{K}}$ of the original kernel matrix $\bm{K}$ has to be accurate enough to ensure that the diagonal normalization matrices $\tilde{\bm{D}}$ and $\tilde{\bm{Q}}$ have positive diagonal values. If this is not the case, the normalization is not well defined and may lead to unpredictable numerical errors. To prevent that, one has to choose a sufficiently large rank parameter $r$. This issue does not arise in the subsampling method because the subsampled matrix $\tilde{\bm{K}}$ is itself a kernel matrix, which will always lead to positive diagonal entries for the corresponding matrices $\tilde{\bm{D}}$ and $\tilde{\bm{Q}}$.
On the other hand, the dilution method has the advantage of incorporating at least partial information about the whole dataset $X_{N}$. This is because the partial Cholesky factorization of $\bm{K}$ involves projecting the matrix onto the linear span of its $r$ selected columns. Thinking of each column $i\in\{0,\ldots,N-1\}$ as corresponding to state $x_{i}\in X_{N}$, every column $i$ sampled by the pivoted partial Cholesky algorithm contains the pairwise distances $k(x_{i},x_{j})$ between state $x_{i}$ and all other states $x_{j}\in X_{N}$, $j\in\{0,\ldots,N-1\}$. For the choice of sampling strategy introduced in [11], this is also reflected in the error bound (1.1) satisfied by the dilution method’s approximation $\tilde{\bm{K}}$. On the contrary, the subsampling method does not utilize any information about the states rejected during the sampling process, and employs a kernel approximation $\tilde{\bm{K}}$ that is not guaranteed to satisfy the appropriate error bound (1.1). In Section 4.2 we demonstrate empirically two ways in which these differences between the two methods may manifest themselves in the obtained eigenvalues.
Finally, for the specific case of the bistochastic normalized kernels considered in this work, the subsampling method has an asymptotic cost $O(Nr^{3})$ which is worse than dilution’s $O(Nr^{2})$. This difference in cost arises in the Nyström extension step due to the nonlocal formula for the point evaluation of the bistochastic kernel $p$. At the same time though, the extension of each eigenfunction to out of sample points is an inherently parallel computation, requiring no interprocess communication for its numerical implementation. The interplay between parallelization and the difference in asymptotic cost will become important for large datasets.
4 Application
We apply the two algorithms for the approximate computation of the EVD of bistochastic normalized kernel matrices to the extraction of patterns from spatiotemporal dynamics. As the dynamical model we consider the Kuramoto-Sivashinsky (KS) equation
$\partial_{t}u=-u\partial_{s}u-\partial_{s}^{2}u-\partial_{s}^{4}u\qquad t\geq 0,\quad s\in S$ (4.1)
with periodic boundary conditions on the one dimensional compact domain S=[L/2,L/2]S=[-L/2,L/2], $L>0$. In the above $u$ denotes the real valued state variable $u(t,\cdot)\in X$, $t\geq 0$ with state space $X\subset H_{S}=L^{2}(S,\nu;\mathbb{R})$ and $\nu$ the Lebesgue measure on $\mathbb{R}$.
The KS equation is a dissipative partial differential equation that displays spatiotemporal chaotic dynamics and was originally introduced as a model for dissipative wave propagation [34, 42]. The bifurcation parameter controlling the complexity of the dynamics is the domain length $L$, with dynamics ranging from steady solutions and traveling waves for low values of $L$ all the way to spatiotemporal chaos for larger values [32, 16]. The transition to chaos takes place through a sequence of period doubling bifurcations [39].
In addition to its rich dynamics and well understood bifurcation diagram, the KS problem (4.1) has additional desirable properties that make it an excellent testbed for pattern extraction methods. First, its generated solutions satisfy symmetries acting on the spatial domain; more specifically, the KS problem (4.1) is equivariant under translations $u(t,x)\mapsto u(t,x+y)$ for all $y\in\mathbb{R}$ and reflection $u(t,x)\mapsto-u(t,-x)$. Second, it has a global attractor of finite dimension; namely, a finite dimensional subset $A\subset X$ which is forward invariant under the dynamics and attracts all initial conditions $u(0,\cdot)\in X\subset H_{s}$ [43, 40].
The dynamics generated by the KS problem (4.1) is given by the flow map $\Phi^{t}:X\to X$ with continuous time variable $t\geq 0$, state space $X$ and invariant probability measure $\mu$ with compact support $\operatorname{supp}\mu\subseteq A$. In what follows we also employ the discrete time flow map $\Phi^{n\Delta t}:X\to X$ with sampling timestep $\Delta t\geq 0$ and $n\in\mathbb{N}$. The real valued observables of the dynamics are members of the Hilbert space of functions $H_{X}=L^{2}(X,\mu;\mathbb{R})$.
4.1 Spatiotemporal pattern extraction
The problem of identifying spatiotemporal patterns of a dynamical system has traditionally been formulated as an eigendecomposition problem for a kernel integral operator $K$ acting on the space of observables $H_{X}$. The most popular method is arguably proper orthogonal decomposition (POD), where the kernel integral operator is formed using a two-point spatial correlation kernel function [2, 6, 29]. In this work we employ an alternative but related approach called vector valued spectral analysis (VSA) [26].
The VSA method employs the product state space $\Omega=X\times S$ and associated product Hilbert space $H_{\Omega}=L^{2}(\Omega,\sigma;\mathbb{R})$ with product measure $\sigma=\mu\times\nu$. As a Hilbert space, $H_{\Omega}$ is isomorphic to the space of vector valued observables $H=L^{2}(X,\mu;H_{S})$ and to the tensor product space $H_{X}\otimes H_{S}$. Every function $f\in H_{\Omega}$ represents a spatiotemporal pattern of the dynamics, with a temporal dependence through $x\in X$ and a spatial dependence through $s\in S$. More specifically, for every $x\in X$ $f(x,\cdot)\in H_{S}$ denotes a function on the spatial domain $S$ with $f(x,s)\in\mathbb{R}$ its pointwise value at $s\in S$. The map $t\mapsto f(\Phi^{t}(x),\cdot)\in H_{S}$ represents the temporal evolution of pattern $f\in H_{\Omega}$ by the dynamics $\Phi^{t}$ for initial state $x\in X$.
The desired spatiotemporal patterns are given by the eigenfunctions of a kernel integral operator $K\colon H_{\Omega}\to H_{\Omega}$
$Kf(\omega)=\int_{\Omega}\kappa(\omega,\omega^{\prime})f(\omega^{\prime})d\sigma(\omega^{\prime})$ (4.2)
with product state $\omega=(x,s)\in\Omega$ and continuous, bounded and positive definite kernel function $\kappa\colon\Omega\times\Omega\to\mathbb{R}$. The integral operator $K$ is compact and selfadjoint; as a result, its eigenfunctions form an orthonormal basis of $H_{\Omega}$ and its eigenvalues are real, nonnegative and have zero as their limit point. By forming an operator acting directly on $H_{\Omega}$ we obtain spatiotemporal patterns that are generally not of tensor product form, meaning that they are not expressible as the tensor product of a pair of temporal and spatial modes. This is in contrast to traditional approaches such as POD, where one computes the eigenfunctions of an operator acting on the temporal space $H_{X}$ and forms their tensor product with a basis for the spatial space $H_{S}$.
For our kernel function $\kappa=k\circ(W_{Q}\times W_{Q})$ we employ a gaussian kernel function $k\colon\mathbb{R}^{Q}\times\mathbb{R}^{Q}\to\mathbb{R}$ acting by composition with a delay embedding map $W_{Q}\colon\Omega\to\mathbb{R}^{Q}$ where $Q\in\mathbb{N}$ denotes the number of time delays. Given a product state sample $\omega=(x,s)\in\Omega$ the delay embedding map $W_{Q}$ forms $Q$ delays in time at the spatial point $s\in S$
$W_{Q}((x,s))=(x(s),\Phi^{-\Delta t}(x)(s),\ldots,\Phi^{-(Q-1)\Delta t}(x)(s))$
and the gaussian kernel function $k$ acts by
$k(W_{Q}(\omega),W_{Q}(\omega^{\prime})=\exp\bigl(-\frac{1}{\epsilon Q}\lVert W_{Q}(\omega)-W_{Q}(\omega^{\prime})\rVert_{2}^{2}\bigr)$ (4.3)
where $\epsilon>0$ is a tunable bandwidth parameter and $\lVert\cdot\rVert_{Q}$ denotes the 2-norm in $\mathbb{R}^{Q}$. For our numerical experiments (Section 4.2) we are going to use the bistochastic normalized versions of kernel (4.3) and associated kernel integral operator (4.2).
Figure 1: (Left) Space-time heatmap of the true state data obtained by integrating the KS problem (4.1) for $500$ time units using the parameter values given in Section 4.2. (Right) Space-time heatmap of the same state data as in the left plot with black dots indicating the states sampled by the pivoted partial Cholesky algorithm for rank parameter $r=2048$.
Thanks to its acting by composition with the delay embedding map $W_{Q}$, kernel function $\kappa$ factors the product state space $\Omega$ into equivalence classes consisting of states with identical dynamical behavior under $Q$ delays. As shown in [26], this implies that the functions in the range $\operatorname{ran}K$ of integral operator (4.2) are invariant under the actions of the symmetry group of the KS problem (4.1). To make this property precise, we consider the group of symmetries $G$ with continuous left action on the spatial domain $\Gamma_{S,g}\colon S\to S$ for $g\in G$. For the KS problem (4.1), the actions $\Gamma_{S,g}$ represent spatial translations and reflection. Every induced action $\Gamma_{X,g}\colon X\to X$, $\Gamma_{X,g}(x)=x\circ\Gamma_{S,g}^{-1}$ represents a dynamical symmetry of the dynamics generated by (4.1), meaning that the dynamics $\Phi^{t}$ satisfies the equivariance property
$\Gamma_{X,g}\circ\Phi^{t}=\Phi^{t}\circ\Gamma_{X,g}$
for all $g\in G$ and $t\geq 0$. For our choice of kernel function $\kappa$, every function $f\in\operatorname{ran}K$ satisfies the analogous invariance property
$f\circ\Gamma_{\Omega,g}=f$ (4.4)
for all $g\in G$ with induced action $\Gamma_{\Omega,g}\colon\Omega\to\Omega$, $\Gamma_{\Omega,g}=\Gamma_{X,g}\otimes\Gamma_{S,g}$. Thanks to the invariance property (4.4), every eigenfunction of the kernel integral operator (4.2) is invariant under the actions of $G$ on $\Omega$, meaning that each such function can generally represent a more complex spatiotemporal pattern than when symmetry invariance is not ensured [26].
Figure 2: Comparison of eigenfunctions $\phi_{1}$ (top row) and $\phi_{2}$ (bottom row) obtained using the true EVD of the bistochastic kernel matrix (leftmost column), the dilution method (middle column) and the subsampling method (rightmost column).
4.2 Numerical experiments
For our numerical experiments we consider the KS problem (4.1) for the domain length value $L=22$ which corresponds to chaotic dynamics. We perform a Fourier spatial discretization using $M=64$ Fourier modes and 3/2 dealiasing of the pseudospectral treatment of the quadratic nonlinearity [9]. For the temporal discretization we employ the exponential time differencing 4-stage Runge-Kutta (ETDRK4) method introduced in [15] and further developed in [31], using the timestep value $\Delta t=0.25$. Our initial condition is formed by setting the leading four Fourier coefficients to $0.6$ and the rest to zero. After integrating for 10 00010\,000 timesteps (2 5002\,500 time units) we collect one sample every four timesteps (one time unit) for a total of $563$ samples. A space-time plot of the integrated solution is shown in the left panel of Figure 1.
Figure 3: Comparison of eigenfunctions $\phi_{4}$ (top row) and $\phi_{5}$ (bottom row) obtained using the true EVD of the bistochastic kernel matrix (leftmost column), the dilution method (middle column) and the subsampling method (rightmost column).
Using $Q=64$ delays, our training dataset consists of $N=500$ samples in delay embedded form, bringing the total number of product state samples to NM=32 500NM=32\,500. To employ the dilution and subsampling algorithms developed in the previous section we must first calibrate the bandwidth parameter $\epsilon$ of kernel (4.3). We perform the bandwidth calibration in two stages. First, we use the median rule studied in [24] and the full training dataset to get a first approximation of parameter $\epsilon$. Second, using the obtained $\epsilon$ value we apply the pivoted partial Cholesky algorithm developed in [11] to sample $8192$ states from our training dataset. Relying only on the sampled subset of training data, we use the scaling-based algorithm developed in [14] to refine our approximation of bandwidth $\epsilon$. The bandwidth value selected by this two-stage algorithm is $\epsilon=15$. After experimenting with similar bandwidth values, the numerical results presented below are obtained using the adjusted value $\epsilon=50$.
Figure 4: Comparison of eigenfunctions $\phi_{3}$ (top row) and $\phi_{6}$ (bottom row) obtained using the true EVD of the bistochastic kernel matrix (leftmost column), the dilution method (middle column) and the subsampling method (rightmost column).
Next we employ the dilution and subsampling algorithms to approximate the EVD of the kernel matrix corresponding to the bistochastic normalized version of kernel (4.3). For that we perform a partial Cholesky factorization with rank parameter $r=2048$, yielding a trace norm error of $7.15\%$. The right panel of Figure 1 shows the product states sampled by the pivoted partial Cholesky algorithm for the used rank parameter.
The dilution and subsampling results are compared with the true results obtained by directly computing the EVD of the bistochastic kernel matrix. More specifically, we test the performance of each method by focusing on two aspects of the approximate results. First, we compare the individual eigenvalues and eigenfunctions obtained by each method with the corresponding true results. Computing an accurate approximation of individual eigenvalues and functions is important in applications such as kernel smoothing, where one uses a kernel integral operator to regularize a function or the spectrum of a linear operator. Second, we test whether the linear span of the eigenfunctions computed by each method is of sufficiently high rank to accurately represent the true state data. This property is important in applications such as reduced modeling and dynamical closure, where one uses a data driven basis to build a compressed representation of observables of the true dynamics.
Figures 2, 3 and 4 compare a selection of the eigenfunctions obtained by the dilution and subsampling algorithms for rank parameter $r=2048$ with the corresponding true eigenfunctions. More specifically, Figure 2 compares eigenfunctions $\phi_{1}$ and $\phi_{2}$, Figure 3 $\phi_{4}$ and $\phi_{5}$, and Figure 4 $\phi_{3}$ and $\phi_{6}$. Since we are using an ergodic bistochastic kernel integral operator, the leading eigenfunction $\phi_{0}$ obtained by both methods is a constant function, which is why we do not include it in our comparisons. The associated eigenvalues obtained by each method are shown in Figure 5; in analogy to $\phi_{0}$, the leading eigenvalue $\lambda_{0}$ obtained by each method is equal to one.
Figure 5: Comparison of the true eigenvalues with those obtained using the dilution and subsampling methods. In the left plot the horizontal axis is limited to the maximum value of 20 00020\,000 to facilitate the visual comparison of the dilution and subsampling eigenvalues. The right plot is a subset of the left one focusing on the leading 20 eigenvalues.
Looking at Figures 2 and 3 we see that the dilution and subsampling methods extract very similar spatiotemporal patterns as the eigenfunctions 1, 2, 4 and 5. In addition, the patterns are in qualitative and quantitative agreement with the true eigenfunctions. Given that these are some of the eigenfunctions corresponding to the largest eigenvalues, it is reasonable to expect that both methods will agree in their results, assuming sufficient sampling of the underlying dynamics.
Comparing eigenfunctions 3 and 6 shown in Figure 4, we see that the two methods extract similar patterns but in exchanged order; namely, the dilution eigenfunctions $\phi_{3}$ and $\phi_{6}$ are respectively similar to the subsampling eigenfunctions $\phi_{6}$ and $\phi_{3}$. In particular, the subsampling eigenfunction $\phi_{6}$ seems to combine features of both $\phi_{3}$ and $\phi_{6}$ obtained by dilution. One explanation for this behavior can be offered by looking at the eigenvalues comparison in Figure 5. There we see that the subsampling eigenvalues $\lambda_{3}$ to $\lambda_{7}$ are very close in magnitude, which means that a small perturbation in their values can reshuffle the order of the patterns extracted as the respective eigenfunctions. In addition, these subsampling eigenvalues are very close to the dilution eigenvalue $\lambda_{3}$, which may be used to explain why the subsampling eigenfunction $\phi_{6}$ seems to incorporate features of the dilution eigenfunction $\phi_{3}$. Overall, the linear span of the leading seven eigenfunctions (including $\phi_{0}$) is very similar between the two methods and in close agreement with the true results.
Figure 6: Comparison of the true state data (leftmost column) for $500$ time units with the state data projected onto the linear span of the dilution eigenfunctions (middle column) and the subsampling eigenfunctions (rightmost column).
Returning to the eigenvalues comparison in Figure 5, one noticeable difference between the two sets of eigenvalues is their rate of decay as the eigenvalue index is growing. In particular, the subsampling eigenvalues have a tail that decays more quickly than the dilution eigenvalues. This is reasonable considering that the subsampling method uses only a subset of the original samples and does not contain any information about the rejected samples. On the contrary, the dilution method employs at least partial information about all training samples (Section 3.3). In this case, this translates to a wider eigenvalue spectrum for the dilution eigenvalues; namely, a slower rate of decay of the magnitude of the eigenvalues which is closer to the true decay rate.
In addition, it appears that the subsampling method overestimates the magnitude of the eigenvalues for moderate indices, compared to the dilution and true results. Combined with the faster decay rate for large indices, this can be interpreted as the subsampling method placing a larger relative importance to the eigenfunctions corresponding to low and moderate indices. In this line of thinking, the relative “importance” of an eigenfunction is defined as the ratio of its eigenvalue over the sum of all eigenvalues, qualitatively similar to the “energy ratio” used in proper orthogonal decomposition [2].
The above empirical observations are backed by the fact that the rank-$r$ kernel approximation $\tilde{\bm{K}}$ employed by the dilution method satisfies the error bound (1.1) for appropriate choices of parameters $r^{\prime}\leq r$ and $\epsilon$, which is generally not true for the subsampling method. This implies that the dilution method’s $\tilde{\bm{K}}$ is close to a spectrally optimal approximation of $\bm{K}$ of rank $r^{\prime}\leq r$, which is reflected in the very close agreement between the dilution and true eigenvalues and eigenfunctions for low to moderate indices.
Figure 7: Comparison of the true state data (leftmost column) for 2 5002\,500 time units with the state data projected onto the linear span of the dilution eigenfunctions (middle column) and the subsampling eigenfunctions (rightmost column).
Next we test whether the linear span of the computed eigenfunctions is of sufficiently high rank to represent the state variables of the underlying spatiotemporal dynamics. To do so, we project the training state data onto the linear span of the eigenfunctions obtained by each method. The projection results are shown in Figure 6, where the two methods are compared with the true state data. The visual comparison indicates that both sets of eigenfunctions can represent the true state data with high accuracy in a pointwise sense. More specifically, the relative 2-norm of the pointwise error is equal to $2.32\%$ for dilution and $2.40\%$ for subsampling, normalized by the 2-norm of the true state data.
To facilitate the comparison of the dilution and subsampling results with the true ones, we have so far restricted ourselves to a training dataset of modest size. We now increase the size of the training dataset by five times to include a total of 2 5002\,500 time units (NM=160  [^0] NM=160\,000) while keeping the approximation rank $r=2048$ and all other parameters the same, simulating a scenario where more aggressive approximation is required. The relative trace norm error of the pivoted partial Cholesky algorithm obtained for this training dataset is $13.25\%$. To test the performance of the two methods for this choice of parameters we focus on the projection of the true state data onto the linear span of the eigenfunctions obtained by each method. The projection results are shown in Figure 7, where we see that both methods can again represent the true state data reasonably well. In particular, the relative 2-norm of the pointwise error is equal to $3.60\%$ for the dilution method and $3.70\%$ for subsampling.
4.3 Discussion
Based on our comparison of the eigenvalues and eigenfunctions obtained by each of the dilution and subsampling methods, we conclude that the dilution method can be expected to yield a more accurate approximation for a given approximation rank. This is reflected in the closer agreement of the dilution eigenvalues and eigenfunctions with the true ones, as well as in the approximation error bound (1.1) satisfied by the low rank approximation $\tilde{\bm{K}}$ employed by the dilution method.
In terms of representing the training state data, the two methods performed similarly for both sizes of the training dataset used in this work. Namely, although there were differences in their individual eigenvalues and eigenfunctions, the linear span of the eigenfunctions obtained by each method was of sufficiently high rank to represent the training state data.
The choice of kernel function used in defining the kernel integral operator (4.2) plays an important role in determining the rank of its range, and ultimately the maximum rank that can be obtained by the linear span of the numerically computed eigenfunctions. Our use of a gaussian kernel means that the rank of the integral operator is infinite, regardless of the rank of the observation map used to generate the employed state data [41, 26]. On the contrary, the state correlation kernels traditionally used in POD do not share that property, meaning that the rank of the associated integral operator is bounded by the rank of the observation map. Although the rank obtained by POD is sufficient to represent the state data to any given level of accuracy, it might not be sufficient to represent arbitrary observables. As a result, our choice of the gaussian kernel (4.3) is motivated by applications where representing observables based on the computed eigenfunctions is important [22, 23, 46].
Finally, for our numerical results we used the gaussian kernel (4.3) with a fixed bandwidth parameter $\epsilon$. The methods considered in this work can be applied without change to kernels of variable bandwidth, where the employed bandwidth depends on the arguments of the kernel function [7, 25].
5 Conclusion
We developed an algorithm for the approximate computation of the eigenvalue decomposition of bistochastic normalized kernel matrices. The proposed algorithm employs a pivoted partial Cholesky algorithm to construct a low rank approximation of the original kernel matrix and compute the approximate eigenvalue decomposition of its bistochastic normalization, relying on a limited number of kernel evaluations. Additionally, we compared the developed algorithm with an alternative based on subsampling and Nyström extension. We applied both algorithms to the kernel based extraction of spatiotemporal patterns from chaotic dynamics, demonstrating their relative performance and investigating their differences.
Next steps in this line of research involve using our proposed algorithm to enable the application of kernel methods to large datasets for tasks such as spatiotemporal pattern extraction, model reduction and dynamical closure.
Acknowledgments
DG acknowledges support from the US Department of Energy under grant DE-SC0025101. CV was supported as a postdoctoral researcher from this grant.